The Sir Nigel Gresley Locomotive Trust Ltd

OVERHAUL 2015 - 2020

All the overhaul photos may be found here.

Sir Nigel Gresley is now out of traffic for overhaul. This will probably take four years and cost about £750,000. If you would like to make a donation towards this overhaul please click on one of the the donate buttons.

To donate via PayPal (incl. credit or debit card):Donate         To donate by 'phone: Donate

Notes from information supplied by Darrin Crone, Locomotive Engineer.

Clicking on the images will display a larger version.
Clicking again will close the window. Where an 'X' shows bottom right of the image, click on this (or use the 'f' key) to display full size.
(You may wish to maximise your screen for this)

From July 2019.

26 July 2019

The welded return cranks were collected from contractors and have been returned to York. The crank angle set up jig has been used to mark the dummy cranks. We were very pleased with how well the jig performed. The dummy cranks which will provide patterns for machining the real cranks have now been taken to be machined. There's a considerable amount of machining work to be done on the cranks and the pressure is on to maintain the quality and accuracy so far achieved.

The tender spring hangers are now being painted while the tender frames are receiving black gloss. The tender spring hanger Spenser rubber boxes and the bottom plates are also now being painted.

The repaired tender spring hangers were returned with NDT documentation covering the repaired areas. That left the hangers not requiring repair, and the hook ends of all of them to NDT. So a last round of examinations were carried out, and a crack was found in one of them. The defect was ground out, as often surface defects can be removed on components like these that show a lot of surface history, however this was definitely a crack and extended in to the hanger. It will require repair.

All of the new tender hornstay bolts have been completed with the drilling of the split pin holes.

The tender brasses continue to be scraped on to the journals. Unfortunately two have had to be reworked by contractors as they had excessive end float on the journal.

The tender front progresses with the lower 2 front plates trimmed, drilled and countersunk, while work has started grinding back the long weld that secured the streamlining to the original 1928 tender top. The exposed frame work at the front has been needle gunned to remove as much of the corrosion as possible.

The butt strap used to increase the height of the front of the tender when streamlined will serve no purpose on the reconstructed tender but will be refitted. The strap was removed complete with all the rivet heads and these have been secured in the strap by welding.

In examining the front of the tender it was noticed that the corridor door appears to have moved and was not closing correctly. We know the tank doesn't quite sit flat so to make sure we don't build a twist in to it the beams the tank sits on were re-levelled. The packing the tank sits on was also rearranged to support the tank along the frame line. This didn't seem to affect the position of the door much, so a string was run along the sides of the tank and it was noticed that the tank side definitely bends in toward the front of the right side. A prop was then put in and jacked out until the door lined up with the exposed frame work in the tank. This required the cutting of the beam put to stabilise the front, which was put in before the front plate work was removed. The plates will be fitted to the current shape of the tank but final securing will take place after the tank is fastened down on to the frames, as if we make the front of the tank too rigid now we could end up deforming the frames.

A template has been made of the top section of the tender tank front, and was used to mark the location of the fixing holes. After the prop was put in the template showed that the holes had shifted. The tank being much more rigid on the side opposite the corridor.

The piping team has been working on running the large diameter air supply pipe and putting in pipe clamps and supports to other pipe runs. Also in the tender frames some final work has been done to the TPWS conduit runs. The OTMR sensor enclosure was weld repaired and has now been drilled for the sensor bulkhead connection holes. The sensor bulkhead fittings have been cleared of contamination and the sensors have been checked. All sensors seem to be doing what they should but they will required formal calibration before refitting.

The middle tender vacuum cylinder bracket was positioned and the fixing holes were drilled through. The bracket has now been removed and the holes are being machined for the screws that go through the bracket and hold down the vacuum cylinder trunions. New screws have also been machined.

The lower vestibule support rod housings are being welded to restore thickness lost to wear.

The first of the loco brake hangers has been rebushed and was clamped to the borer table to examine its straightness, as when on the loco some of the hangers appeared to be twisted. On the table compared to a large square there was some misalignment between the end bushes but it was very small, probably within tolerance. Attention then turned to the trailing pins in the loco frames. These are harder to get at than the others and do show more wear, probably being older. Due to the load on the pin they have worn out of shape and this may be the cause of the apparent twist seen when the loco came out of traffic. It has been decided to replace these pins as they are too worn to be trued up and reused.

Work continues on the valve gear with the finishing the fitting of the pin in the reverser column crank. At the other end of the rod the lifting arms and nuts have been cleaned up by dressing out as much of the surface damage as we can without excessive removal of material, or character. The leading pins on the radius rods have also now been fitted and he is now on to the combination levers.

Work on the slidebar bolts continues with the right side being worked on.

The big end bearing pattern has now been received and was promptly sent to the foundry. They are a little concerned that the pattern won't give us enough material on the bore to machine to the dimensions we need, so additional work will be required on the pattern before casting.

The strip down, clean and examination of the cods mouth door operating gear continues.

At Llangollen the superheater header is now successfully fitted and has been hydraulically tested in position. This means work can now commence on fitting the superheater elements.

A full set of new tender hornstay bolts. 13 July 2019
Photograph: Darrin Crone

Tender hornstay bolts

The tender frames are now painted to black gloss. 13 July 2019
Photograph: Darrin Crone

Tender frames painted

The valve gear lifting arms, not at the correct angle, have now been overhauled to accept new pins. 13 July 2019
Photograph: Darrin Crone

Valve gear lifting arms

New pins have also been fitted to the valve gear radius rods. 13 July 2019
Photograph: Darrin Crone

New pins have been fitted

The superheater header in place and being pressure tested on 16 July 2019
Photograph: Paul Aston

The superheater header

A pattern has been made of the plates required for the top of the tender front. 17 July 2019
Photograph: Peter Bracckstone

Top of the tender front

The jig used to accurately identify the angle that the return cranks should be set. 18 July 2019
Photograph: Richard Swales

Return cranks jig

On the back of the dummy return cranks the distance from the wheel centre is marked and the length of the crank from the centre of the pin. 18 July 2019
Photograph: Richard Swales

Dummy return cranks

The marked dummy return cranks will be used as patterns for the machining of the real return cranks to the correct dimensions. 18 July 2019
Photograph: Richard Swales

Marked dummy cranks

12 July 2019

The coach has been visited by our CME and myself during the last week in June. We were impressed by the quality of the work we saw at CTMS, who are the contractors carrying out the work. One side of the coach has been extensively re-skinned while the other side has been stripped back to metal. This gives a striking before and after view. It has also shown how much more extensive is the work required to the coach skin and structure.

Meanwhile at Llangollen progress has slowed due to difficulties with sealing the superheater header. It is essential that the header makes a good seal on to the front tubeplate as when the elements are fitted they are only removable with a gas axe. So, the header has been plugged and mounted in the smokebox, and hydraulic-ed. Under test a leak was found. The joint is now being worked on to ensure a reliable seal is achieved. The job is made more awkward as the smokebox is in place and the header has to be angled to be put in or removed. A job made more difficult by all the plugs protruding from the header. The smokebox also makes accessing the joint more difficul.

The tender bearings have now all been remetalled and machined at contractors. The bearings are now at York and are being scraped to their journals. LNER drawings show the back of the tender axleboxes were sealed against the ingress of contamination with a rubber formed seal that was clamped under a plate on the back of the axlebox and ran on the axle near the face of the wheel. These are absent from our axleboxes. We plan to replace a form of sealing, though made from flat sheet. Measurements of the required seal have been taken and quotes for the material will now be obtained.

The last of the tender spring hanger bolts have now been collected from contractors where they have been refurbished.

The steam heat valve that failed during the boiler test at Llangollen, and the patterns for a new valve body, and the new valve guide have now been sent to the foundry.

The tender pipework continues to make progress with a way through found the tender frames for the main air supply run from the air pump to the tender top tanks. This pipe is larger than the other air pipes so this has been a challenge and will require some difficult forming. The pipe previously ran through the cab and along the corridor. Elsewhere the OTMR sensors have been removed from the sensor box and it has been stripped back to metal. The box gland plate was removed, a new seal made and has been sealed in place.

The cab pipework has now been removed leaving the cab floor clear. This has allowed the floor to be finish painted. The reverser stand has also been removed so that there is no chance of damage during the fitting of the boiler. The pipework will also be left out until after the fitting of the boiler to prevent possible damage. So the back of the loco is looking like it did a couple of years ago.

The new cotters for the tender vestibule connection support rods have now been finished. The bolts used to connect the rods to the end face of the coach connection have been machined and the old nuts will be reused.

All of the new tender hornstay bolts have been machined to correct diameter, lathe threaded and the ends parted to length. They only now require drilling for split pins. The tender frames are now receiving black gloss.

All the new bushes for the lower tender pull rods have been manufactured and a number now fitted. The tender handbrake refurbishment is now being specified. The middle tender vacuum brake cylinder bracket has been machined and is temporarily mounted in position. Work has begun on accurately aligning it with the outer brackets and then the mounting holes will be drilled through.

The loco brake hangers have previously had their old bushes removed and the worn faces where the blocks are mounted rebuild with weld. New material has now been obtained for rebushing the large bottom holes. The trailing right hanger has been put up on the loco pin and it is being examined for deformation.

The GSMR enclosure has now been painted to blue gloss as part of the electrical systems overhaul. The new TPWS junction box at the front of the tender frames has now been fitted.

Work continues on the fitting of the new tender front platework. New steel has now been delivered to York. The front section being replaced is formed of four plates and two of these are now in position and being trimmed and drilled for fastening to existing structure. The staging at the rear of the tender has now been moved to the right hand side of the tank to allow the grinding off, of the remains of the streamlining where it was welded to the top of the original 1928 tender.

The drop grate operating mechanism components have been retrieved from store. They were given a clean, fitted in place and examined. There is a plain bronze thrust bearing that will have to be replaced as it is worn completely through.

The cods mouth operating mechanism has now been moved in to the open, and is now stood on trestles in front of the loco. It is now being cleaned and examined.

New pins have now been fitted to the lifting arms and the die blocks have been machined to accept the new pins. The bearing faces in the lifting arms required a little easing and this has now been completed.

The fitting of the new slidebar bolts continues with all the hole gauging now complete. As all the final bolt sizes are now known they are being machined and fitted.

A new, reproduction builders plate has been donated by Mr Suter of York, who handed it over personally at York. It looks superb and it is planned to mount it on the driver's cab side. Anybody wanting to donate a new fireman's side plate?

The cab floor has been stripped out to allow the structure to be finish painted. 25 June 2019
Photograph: Peter Brackstone

The cab floor

Pallets loaded with tender springs return from refurbishment on 25 June 2019
Photograph: Peter Brackstone


The coach stripped back to metal prior to skin repairs at the workshops of CTMS. 26 June 2019
Photograph: Richard Swales

Coach stripped to metal

The coach side that has been repaired. 26 June 2019
Photograph: Richard Swales

Repaired coach side

.Components from the drop grate operating mechanism awaiting painting on 3 July 2019
Photograph: Peter Brackstone

Drop grate components
Return to top

From January to March 2016.
From April to June 2016.
From July to September 2016.
From October to December 2016
From January to March 2017.
From April to June 2017.
From July to September 2017.
From October to December 2017.
From January to March 2018.
From April to June 2018.
From July to September 2018.
From October to December 2018.
From January to March 2019.
From April to June 2019.

Copyright © The Sir Nigel Gresley Locomotive Trust Ltd